# Establishing Linear Surrogate Regret Bounds for Convex Smooth Losses via Convolutional Fenchel–Young Losses

### v2

@@ 3.3 Linear Surrogate Regret Bounds
Consider a target loss with $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For an negentropy function $\Omega:\mathbb{R}^{\varrho}\to\overline{\mathbb{R}}$, suppose Condition 1 holds. For any $\bm{\pi}$-argmax link $\varphi$, $(L_{\Omega_{T}},\varphi)$ admits the following surrogate regret bound:
$\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})\leq N\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta}),~{}~{}~{}\text{for any~{}}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K}.$ (14)
The regret bound (13) indeed has the linear rate $\psi(r)=Nr$ in the form (1), while $L_{\Omega_{T}}$ can be naturally made convex and smooth as discussed in Section 3.2. This is a remarkable consequence because many authors have previously implied the impossibility to overcome the barrier of the square-root regret rate with convex smooth surrogate losses [33, 53, 25, 4]. The crux of this success lies in the infimal convolution structure in (10), which enables us to additively decompose the conjugate $\Omega_{T}^{*}$, exploited below. The constant $N$ will be improved in Section 3.4.
@@
Assume the same set of conditions as in Theorem 13. For any $\bm{\theta}\in\mathbb{R}^{\varrho}$ and $\bm{\pi}\in\Pi(\bm{\theta})$, we have the following inequality:
$\sum_{t=1}^{N}\pi_{t}\mathtt{Regret}_{\ell}(t,\bm{\eta})\leq\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta}),~{}~{}~{}\text{for any~{}}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K}.$ (15)
Its proof, formally given in Appendix A, hinges on the following additive regret decomposition:
$\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta})=\underbrace{R_{L_{\Omega}}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi},\bm{\eta})}_{\text{Risk of Fenchel--Young loss $L_{\Omega}$ $\geq 0$}}+\underbrace{\sum\nolimits_{t\in\widehat{\mathcal{Y}}}\pi_{t}\mathtt{Regret}_{\ell}(t,\bm{\eta})}_{\text{Convex combination of the target regret}}\!\forall(\bm{\theta},\bm{\pi})\in\mathbb{R}^{\varrho}\times\Pi(\bm{\theta}),$
which directly implies the lower bound (15) because the Fenchel–Young loss $L_{\Omega}$ is non-negative. Note that this additive decomposition is indispensable for the linear lower bound (15). This becomes possible just because we create the convolutional negentropy $\Omega_{T}$ based on the additive form in (8), which yields the additive form $\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}$ in the conjugate expression (10). All of these are thanks to the structure of the infimal convolution. With this lemma, Theorem 13 can be proven immediately.
Proof of Theorem 13 Fix any $\bm{\theta}\in\mathbb{R}^{\varrho}$ and choose any $\bm{\pi}\in\Delta^{N}$ out of $\Pi(\bm{\theta})$. Then we have
$\geq\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})/N.$ (by definition of $\varphi$ in Defintion 10)
Finally, we achieve the main result by combining Theorems 11 and 13. We constructively prove the existence of convex smooth linear-regret surrogate losses, by noting that both strongly convex $\Omega$ satisfying Condition 1 and $\bm{\pi}$-argmax link do exist.
@@
Consider a target loss $\ell$ with $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For an negentropy function $\Omega:\mathbb{R}^{\varrho}\to\overline{\mathbb{R}}$ satisfying Condition 1, suppose that $\Omega$ is strongly convex on $\mathrm{dom}(\Omega)$. Then the convolutional Fenchel–Young loss $L_{\Omega_{T}}(\cdot,y)$ is convex, differentiable, and smooth over $\mathbb{R}^{\varrho}$ for any $y\in\mathcal{Y}$. Moreover, with $\bm{\pi}$-argmax link $\varphi$, $(L_{\Omega_{T}},\varphi)$ enjoys the linear surrogate regret bound (14).
@@ 3.4 Improving Constant of Linear Surrogate Regret Bounds
The constant $N$ in the linear surrogate regret bound in (14) can be prohibitively large, leading to a vacuous bound. This issue is significant in structured prediction as studied in Osokin et al. [50]. For example, in multilabel classification with the Hamming loss, $N=2^{\varrho}$ is the number of all the potential binary label predictions that increases exponentially with $\varrho$, the number of binary labels. In top-$k$ classification, $N=\binom{K}{k}$ is the number of all possible size-$k$ subsets of the class space. Thankfully, the geometry property of the problem (10) provides a promising scheme for reducing the dependency on prediction number $N$, which induces the following improved link.
@@
Suppose Condition 1 holds. There exists $\tilde{\bm{\pi}}:\mathbb{R}^{\varrho}\to\Delta^{N}$ such that $\tilde{\bm{\pi}}(\bm{\theta})\in\Pi(\bm{\theta})$ and $\|\tilde{\bm{\pi}}(\bm{\theta})\|_{0}\leq\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1$ for any $\bm{\theta}\in\mathbb{R}^{\varrho}$. Moreover, with the induced $\tilde{\bm{\pi}}$-argmax link $\varphi_{\tilde{\bm{\pi}}}:\mathbb{R}^{\varrho}\to\Delta^{N}$, $(L_{\Omega_{T}},\varphi_{\tilde{\bm{\pi}}})$ admits the following surrogate regret bound:
$\mathtt{Regret}_{\ell}(\varphi_{\tilde{\bm{\pi}}}(\bm{\theta}),\bm{\eta})\leq[\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1]\,\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta}),\quad\text{for any~{}}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K},$ (16)
@@
Proof.  Fix an arbitrary $\bm{\theta}\in\mathbb{R}^{\varrho}$. Since $\Pi(\bm{\theta})$ is non-empty, we can select $\bm{\pi}$ from $\Pi(\bm{\theta})$. Note that $\mathcal{L}^{\bm{\rho}}\bm{\pi}\in\mathrm{conv}(\{\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}})$. According to Carathéodory’s theorem, there exists $\tilde{\bm{\pi}}$ such that $\|\tilde{\bm{\pi}}\|_{0}\leq\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1$ and $\mathcal{L}^{\bm{\rho}}\tilde{\bm{\pi}}=\mathcal{L}^{\bm{\rho}}\bm{\pi}$ is also in $\Pi(\bm{\theta})$. Since $\bm{\theta}$ is chosen arbitrarily, we can then construct a single-valued function $\tilde{\bm{\pi}}:\mathbb{R}^{\varrho}\to\Delta^{N}$ such that $\tilde{\bm{\pi}}(\bm{\theta})\in\Pi(\bm{\theta})$ and $\|\tilde{\bm{\pi}}(\bm{\theta})\|_{0}\leq\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1$. Noting that $\max_{t\in\widehat{\mathcal{Y}}}\tilde{\pi}_{t}(\bm{\theta})>{1}/[\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1]$, we can complete the rest of the proof similarly to Theorem 13.
Note that we have $\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})\leq\min\{N,\varrho\}$, which indicates that the dependency on $N$ can be largely reduced when $\varrho\ll N$. For example, in top-$k$ classification, $\bm{\rho}(y)=\bm{e}_{y}$ is used with $\varrho=K$, which is much smaller than $N=\binom{K}{k}$. In multilabel classification with the Hamming loss, $\varrho=\ln_{2}N$ is logarithmically smaller than $N$ when the binary encoding $\bm{\rho}(y)=\bm{\nu}(y)$ is used (see Section 2.1).
@@ 3.5 Bonus: Fisher-consistent Probability Estimator
In this subsection, we discuss a benefit of convex smooth surrogate losses beyond discrete prediction problems. Oftentimes people are interested in a probability estimator over possible prediction outcomes, recovering from surrogate risk minimizers, as in selective classification [16]. Herein non-smooth surrogate losses have been unfavored because of lacking reasonable probability estimators [42, 15]. Ramaswamy et al. [53] conjectures the impossibility to establish a probability estimator and linear regret bound simultaneously. By contrast, we give a Fisher-consistent probability estimator of the linear property $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$ under the convolutional Fenchel–Young loss, without sacrificing the fast regret bound.
@@
Suppose Condition 1 holds and $\Omega$ is strictly convex on $\mathrm{dom}(\Omega)$. For any $\bm{\eta}\in\mathrm{relint}(\Delta^{K})$, the surrogate risk $R_{L_{\Omega_{T}}}(\cdot,\bm{\eta})$ is minimized at $\bm{\theta}^{*}\in\mathbb{R}^{\varrho}$ such that $\nabla\Omega^{*}(\bm{\theta}^{*}+\mathcal{L}^{\bm{\rho}}\bm{\pi})=\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$ for any $\bm{\pi}\in\Pi(\bm{\theta}^{*})$.
According to the result above, we can first solve $\Pi(\bm{\theta})$ in (12) and select $\bm{\pi}\in\Pi(\bm{\theta})$ to use $\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ as an estimate of $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$ with consistency guarantee. When we follow the trivial loss decomposition (3) with $\bm{\rho}(y)=\bm{e}_{y}$, the estimand is $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]=\bm{\eta}\in\Delta^{K}$. For this reason, we say $\nabla\Omega^{*}(\bm{\theta}^{*}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ is a “probability” estimator. For example, let us recap the encoding of multilabel classification in Section 2.1, where $y\in\mathcal{Y}$ is a multilabel among all $2^{\varrho}$ possible combinations of $\varrho$ binary labels. A multilabel $y$ is encoded by $\bm{\nu}(y)\in\{0,1\}^{\varrho}$, and the Hamming loss is decomposed with $\bm{\rho}(y)=1-2\bm{\nu}(y)$. Here $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\nu}(y)]$ reads $\bar{\bm{\nu}}=[\mathrm{Prob}(\bm{\nu}_{i}(y)=1)]_{i=1}^{\varrho}$, whose $i$-th element indicates the likelihood of the binary class $i$ being positive. Through the relation $\bm{\rho}=1-2\bm{\nu}$, the probability estimator $\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ is capable of recovering $\bar{\bm{\nu}}$ by $[1-\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})]/2$.
@@ 4 Example: Multiclass Classification
In multiclass classification, the class space and the prediction space are the same, that is, $\mathcal{Y}=\widehat{\mathcal{Y}}=[K]$. For an input instance with multiclass probability $\bm{\eta}\in\Delta^{K}$, the goal is to predict the most likely class label $t\in\mathrm{argmax}_{t\in\mathcal{Y}}\,\eta_{t}$. The standard target loss is the 0-1 loss $\ell_{01}(t,y)=[\![t\not=y]\!]$.
Firstly, we adopt the decomposition (3) for this task, that is, $\bm{\ell}^{\bm{\rho}}(t)=[\ell_{01}(t,1),\cdots,\ell_{01}(t,K)]=\bm{1}-\bm{e}_{t}$ and $\bm{\rho}(y)=\bm{e}_{y}$, which corresponds to the one-hot encoding commonly used in this task. In this case, $N=\varrho=K$, and $\mathcal{L}^{\bm{\rho}}=\bm{1}\bm{1}^{\top}-I$, where $I$ is the identity matrix.
Next, we move on to the convolutional negentropy $\Omega_{T}$. For the choice of $\Omega$, we use the Shannon negentropy, which induces the celebrated cross-entropy loss in the original Fenchel–Young loss framework [10, Table 1]. Its conjugate is the log-sum-exp function $\Omega^{*}(\bm{\theta})=\ln\langle\exp(\bm{\theta}),\bm{1}\rangle$. Then we can calculate the conjugate of $\Omega_{T}$ based on Lemma 8:
@@
$L_{\Omega_{T}}(\bm{\theta},y)=\Omega_{T}^{*}(\bm{\theta})+\Omega_{T}(\bm{\rho}(y))-\langle\bm{\theta},\bm{\rho}(y)\rangle=\min_{\bm{\pi}\in\Delta^{K}}\ln\left\langle\exp(\bm{\theta}+\bm{1}-\bm{\pi}),\bm{1}\right\rangle-\theta_{i},$ (18)
which is ensured to be convex and smooth in $\bm{\theta}$ thanks to Theorem 11. While sharing the form of cross-entropy loss $L_{\Omega}(\bm{\theta},y)=\ln\langle\exp(\bm{\theta}),\bm{1}\rangle-\theta_{y}$, it further incorporates an additional bounded perturbation.
@@
1:  Sort $\bm{\theta}\in\mathbb{R}^{K}$ such that $\theta_{[1]}\geq\cdots\geq\theta_{[K]}$.
2:  $n\leftarrow\max\{k\in[K]:~{}1+k\theta_{[k]}>\sum_{i=1}^{k}\theta_{[i]}\}$.
3:  $\tau(\bm{\theta})\leftarrow\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}$.
4:  $\pi_{\log}(\theta)_{i}\leftarrow\max\{\theta_{i}-\tau(\bm{\theta}),0\}$.
@@
The proof is deferred to Appendix B.1, following from the KKT conditions. Eventually we have
@@
Algorithm 1 comes with a significant resemblance with the sparsemax computation [41, Algorithm 1], which is determined by the similar structure shared by (17) and Euclidean projection problem [41, (2)]. Since $\Pi(\bm{\theta})$ is a singleton, we have the unique $\bm{\pi}_{\log}$-argmax link $\varphi_{\bm{\pi}_{\log}}$ (Definition 10). While we need to solve the problem (17) for every $\bm{\theta}$ to have access to $\nabla_{\bm{\theta}}L_{\Omega_{T}}(\bm{\theta},y)$, the test-time prediction by the $\bm{\pi}_{\log}$-argmax link is much cheaper, without requiring Algorithm 1. The proof can be found in Appendix B.2.
@@
According to this result, we can simply use the class label with the largest score in test time.
@@
Under classification, a surrogate loss with a regret bound is classification-calibrated, which indicates that the surrogate risk minimization eventually leads to the Bayes-optimal classifier. In literature, Blondel [8] and Wang and Scott [68] have investigated sufficient conditions for Fenchel–Young losses to be classification-calibrated. Therein the base negentropy is assumed to be of Legendre-type or twice differentiable. Our convolutional Fenchel–Young losses are interesting because we do not require these conditions to yield both the smoothness and a regret bound. Thus it remains open to relax these existing sufficient conditions for Fenchel–Young losses further.
5 Discussions

### v3

@@ 3.3 Linear Surrogate Regret Bounds
Consider a target loss with $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For a negentropy $\Omega:\mathbb{R}^{d}\to\overline{\mathbb{R}}$, suppose Condition 1 holds. For any $\bm{\pi}$-argmax link $\varphi$, $(L_{\Omega_{T}},\varphi)$ admits the following surrogate regret bound:
$\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})\leq N\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta}),~~~\text{for any~}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{d}\times\Delta^{K}.$ (14)
The regret bound in Theorem 13 indeed has the linear rate $\psi(r)=Nr$ in the form (1), while $L_{\Omega_{T}}$ can be naturally made convex and smooth as discussed in Section 3.2. This is a remarkable consequence because many authors have previously implied the impossibility to overcome the barrier of the square-root regret rate with convex smooth surrogate losses [mahdavi, rama_BEP, poly_regret, regret_proper2]. The crux of this success lies in the infimal convolution structure in (10), which enables us to additively decompose the conjugate $\Omega_{T}^{*}$, exploited below. The constant $N$ will be improved in Section 3.4.
@@
Assume the same set of conditions as in Theorem 13. For any $\bm{\theta}\in\mathbb{R}^{d}$ and $\bm{\pi}\in\Pi(\bm{\theta})$, we have the following inequality:
$\sum_{t=1}^{N}\pi_{t}\mathtt{Regret}_{\ell}(t,\bm{\eta})\leq\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta}),~~~\text{for any~}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{d}\times\Delta^{K}.$ (15)
Its proof, formally given in Appendix B, hinges on the following additive regret decomposition:
[…]
which directly implies the lower bound (15) because the Fenchel–Young loss $L_{\Omega}$ is non-negative. Note that this additive decomposition is indispensable for the linear lower bound (15). This becomes possible just because we create the convolutional negentropy $\Omega_{T}$ based on the additive form in (8), which yields the additive form $\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}$ in the conjugate expression (10). All of these are thanks to the structure of the infimal convolution, and Theorem 13 immediately follows with rescaling.
Finally, we achieve the main result by combining Corollary 11 and Theorem 13. We constructively prove the existence of convex smooth linear-regret surrogate losses, by noting that both strongly convex $\Omega$ satisfying Condition 1 and $\bm{\pi}$-argmax link do exist.
@@
Consider a target loss $\ell$ with $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For a negentropy $\Omega:\mathbb{R}^{d}\to\overline{\mathbb{R}}$ satisfying Condition 1, suppose that $\Omega$ is strongly convex on $\mathrm{dom}(\Omega)$. Then the convolutional Fenchel–Young loss $L_{\Omega_{T}}(\cdot,y)$ is convex, differentiable, and smooth over $\mathbb{R}^{d}$ for any $y\in\mathcal{Y}$. Moreover, with $\bm{\pi}$-argmax link $\varphi$, $(L_{\Omega_{T}},\varphi)$ enjoys the linear surrogate regret bound (14).
@@ 3.4 Improving Constant of Linear Surrogate Regret Bounds
The constant $N$ in the linear surrogate regret bound in (14) can be prohibitively large, leading to a vacuous bound. This issue is significant in structured prediction as studied in Self_osokin. For example, in multilabel classification with the Hamming loss, $N=2^{d}$ is the number of all the potential binary label predictions that increases exponentially with $d$, the number of binary labels. In top-$k$ classification, $N=\binom{K}{k}$ is the number of all possible size-$k$ subsets of the class space. Thankfully, the geometry property of the problem (10) provides a promising scheme for reducing the dependency on prediction number $N$, which induces the following improved link.
@@
[…]
$\mathtt{Regret}_{\ell}(\varphi_{\tilde{\bm{\pi}}}(\bm{\theta}),\bm{\eta})\leq[\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1]\,\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta}),\quad\text{for any~}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{d}\times\Delta^{K},$ (16)
@@
Note that we have $\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})\leq\min\{N,d\}$, which indicates that the dependency on $N$ can be largely reduced when $d\ll N$. For example, in top-$k$ classification, $\bm{\rho}(y)=\bm{e}_{y}$ is used with $d=K$, which is much smaller than $N=\binom{K}{k}$. In multilabel classification with Hamming loss, $d=\log_{2}N$ is logarithmically smaller than $N$ when binary encoding $\bm{\rho}(y)=\bm{\nu}(y)$ is used (see Section 2.1).
@@ 3.5 Bonus: Fisher-consistent Probability Estimator
We discuss a benefit of convex smooth surrogate losses beyond discrete prediction problems. Oftentimes people are interested in a probability estimator over possible prediction outcomes, recovering from surrogate risk minimizers, as in classification with rejection [chow1970optimum, bart_r, cortes, cortes2, caoyuzhou, mao_reject, jessie_abs]. Herein non-smooth surrogate losses have been unfavored because of lacking reasonable probability estimators [masnadi2008design, Nontawat]. rama_BEP conjectures the incompatibility of probability estimation with linear regret bounds. By contrast, we give a Fisher-consistent estimator of the linear property $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$ for convolutional Fenchel–Young losses without sacrificing the linear regret bound.
@@
Suppose Condition 1 holds and $\Omega$ is strictly convex on $\mathrm{dom}(\Omega)$. For any $\bm{\eta}\in\mathrm{relint}(\Delta^{K})$, the surrogate risk $R_{L_{\Omega_{T}}}(\cdot,\bm{\eta})$ is minimized at $\bm{\theta}^{*}\in\mathbb{R}^{d}$ such that $\nabla\Omega^{*}(\bm{\theta}^{*}+\mathcal{L}^{\bm{\rho}}\bm{\pi})=\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$ for any $\bm{\pi}\in\Pi(\bm{\theta}^{*})$.
As a result, the empirical minimizers w.r.t. $L_{\Omega_{T}}$ are Fisher-consistent estimators of the linear property $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$. According to the result above, we can first solve $\Pi(\bm{\theta})$ in (12) and select $\bm{\pi}\in\Pi(\bm{\theta})$. Then, $\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ is a rational estimate of the linear property $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$. When we follow the trivial loss decomposition (3) with $\bm{\rho}(y)=\bm{e}_{y}$, the estimand is $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]=\bm{\eta}\in\Delta^{K}$. For this reason, we say $\nabla\Omega^{*}(\bm{\theta}^{*}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ is a “probability” estimator.
For example, let us recap the encoding of multilabel classification in Section 2.1, where $y\in\mathcal{Y}$ is a multilabel among all $2^{d}$ possible combinations of $d$ binary labels. A multilabel $y$ is encoded by $\bm{\nu}(y)\in\{0,1\}^{d}$, and the Hamming loss is decomposed with $\bm{\rho}(y)=1-2\bm{\nu}(y)$. Here $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\nu}(y)]$ reads $\bar{\bm{\nu}}=[\mathrm{Prob}(\bm{\nu}_{i}(y)=1)]_{i=1}^{d}$, whose $i$-th element indicates the likelihood of the binary class $i$ being positive. Through the relation $\bm{\rho}=1-2\bm{\nu}$, the probability estimator $\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ is capable of recovering $\bar{\bm{\nu}}$ by $1-\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})]/2$.
@@ 4 Example: Multiclass Classification
We demonstrate convolutional Fenchel–Young losses for multiclass classification. and further examples of prediction problems can be found in Appendices D and E, including detailed computational complexity analyses and visualizations of the associated binary classification losses. In multiclass classification, the class space and the prediction space are the same: $\mathcal{Y}=\widehat{\mathcal{Y}}=[K]$. For an input with class probability $\bm{\eta}\in\Delta^{K}$, the goal is to predict the most likely class $t\in\mathrm{argmax}_{t\in\mathcal{Y}}\,\eta_{t}$. The target loss is the 0-1 loss $\ell_{01}(t,y)=[\![t\not=y]\!]$.
Firstly, we adopt the decomposition (3) for this task, that is, $\bm{\ell}^{\bm{\rho}}(t)=[\ell_{01}(t,1),\cdots,\ell_{01}(t,K)]=\bm{1}-\bm{e}_{t}$ and $\bm{\rho}(y)=\bm{e}_{y}$, which corresponds to the one-hot encoding commonly used in this task. In this case, $N=d=K$, and $\mathcal{L}^{\bm{\rho}}=\bm{1}\bm{1}^{\top}-I$.
Next, we move on to the convolutional negentropy $\Omega_{T}$. For the choice of $\Omega$, we use the Shannon negentropy, which induces the celebrated cross-entropy loss in the original Fenchel–Young loss framework [FY, Table 1]. Its conjugate is the log-sum-exp function $\Omega^{*}(\bm{\theta})=\ln\langle\exp(\bm{\theta}),\bm{1}\rangle$. Then we can calculate the conjugate of $\Omega_{T}$ based on Lemma 8:
@@
$L_{\Omega_{T}}(\bm{\theta},y)=\Omega_{T}^{*}(\bm{\theta})+\Omega_{T}(\bm{\rho}(y))-\langle\bm{\theta},\bm{\rho}(y)\rangle=\min_{\bm{\pi}\in\Delta^{K}}\ln\left\langle\exp(\bm{\theta}+\bm{1}-\bm{\pi}),\bm{1}\right\rangle-\theta_{y},$ (18)
which is convex and smooth in $\bm{\theta}$ thanks to Theorem 11. While shares the form of cross-entropy loss $L_{\Omega}(\bm{\theta},y)=\ln\langle\exp(\bm{\theta}),\bm{1}\rangle-\theta_{y}$, it further incorporates an additional bounded perturbation.
@@
1: Sort $\bm{\theta}\in\mathbb{R}^{K}$ such that $\theta_{(1)}\geq\cdots\geq\theta_{(K)}$.
[…]
@@
The proof is deferred to Appendix C.1, following from the KKT conditions. Eventually we have
@@
Algorithm 1 comes with a significant resemblance with the sparsemax  [msparsemax, Algorithm 1], which is determined by the similar structure shared by (17) and Euclidean projection problem [msparsemax, (2)]. Since $\Pi(\bm{\theta})$ is a singleton, we have the unique $\bm{\pi}_{\log}$-argmax link $\varphi_{\bm{\pi}_{\log}}$ (Definition 10). While we need to solve problem (17) for every $\bm{\theta}$ to have access to $\nabla_{\bm{\theta}}L_{\Omega_{T}}(\bm{\theta},y)$, the prediction by the $\bm{\pi}_{\log}$-argmax link is much cheaper, without requiring Algorithm 1, which indicates that we can simply use the class label with the largest score $\max_{y\in[K]}\theta_{y}$ in test time. The proof can be found in Appendix C.2.
@@
Under classification, a surrogate loss with a regret bound is classification-calibrated, which indicates that the surrogate risk minimization eventually leads to the Bayes-optimal classifier. In literature, Self3 and wang_unified have investigated sufficient conditions for Fenchel–Young losses to be classification-calibrated. Therein the base negentropy is assumed to be of Legendre-type or twice differentiable. Our convolutional Fenchel–Young losses are interesting because we do not require these conditions to yield both the smoothness and a regret bound. Thus it remains open to relax these existing sufficient conditions for Fenchel–Young losses further.
5 Discussion
Convex non-smooth linear regret losses.
Compared to existing convex non-smooth surrogates with linear regret, e.g., polyhedral losses [total_poly], our smooth convex surrogate has several advantages. First, smoothness enables more efficient optimization, as supported by results in both deterministic and stochastic optimization regimes [yuri, SAG] while it is left open to exploit specific structures arising from polyhedral losses to achieve faster optimization. In addition, the smoothness can lead to optimistic ERM rates of estimation error [smooth1], offering better estimation in easier tasks. Another appealing aspect is that our loss admits a consistent probability estimator (Theorem 17), which can be valuable for downstream tasks such as uncertainty quantification and calibration.